### v1

@@
[…]
1 Introduction
Monte Carlo (MC) simulation is a foundational tool in computational finance and other domains where closed-form solutions for complex stochastic models are unavailable. Its principal virtues are unbiasedness and the ability to rigorously quantify uncertainty through confidence intervals. By relying on random sampling, MC provides well-understood statistical guarantees and error assessments. However, Monte Carlo methods face significant challenges when applied to complex, nested, or path-dependent settings. As Glasserman (2013a) notes,“The slow convergence of Monte Carlo - its $O(1/\sqrt{n})$ error after $n$ replications - is a persistent obstacle to its wider use.” In other words, while MC excels at delivering unbiased estimates with reliable uncertainty quantification, it often becomes a “slow evaluation”. For example, in the context of derivative pricing, the challenge is further compounded for path-dependent options. Hull (2018) points out that “for path-dependent options, the entire price path must be simulated, significantly increasing computational complexity.” Similarly, Broadie and Glasserman (1996) state that “in many financial simulations, the value at each time step is a function of the previous step, creating a sequential dependency that hinders straightforward parallelization.”
@@ 1 Introduction
In recent years, machine learning (ML) prediction methods have emerged as a fast alternative for approximating expensive functions, including prices of options and of other financial derivatives (Ferguson and Liang (2018), Bayer and Stemper (2019)). Neural networks and other ML techniques can learn complex mappings quickly, and once trained, can generate predictions at negligible marginal cost - what practitioners often call ”fast evaluation” (Horváth et al. (2021)). However, their black-box nature and lack of inherent statistical control make it difficult to dissect error and impossible to ensure unbiasedness Hutchinson et al. (1994). As a result, pure ML approaches often yield biased estimates with no guaranteed measure of error, making them potentially unsuitable for high-stakes applications like risk management where reliable uncertainty quantification is crucial Buehler et al. (2019).
This tension between MC’s reliability and ML’s efficiency motivates the exploration of classical variance reduction techniques, particularly control variates (CV). Control variates provide a systematic approach to improving MC efficiency by exploiting auxiliary functions with known expectations ( Glasserman (2013b), Asmussen and Glynn (2007)). When one finds a quantity that is highly correlated with the target while also possessing an analytically known mean, then the variance reduction could be substantial (Boyle et al. (1997)). However, as in many path-dependent exotic derivatives, under many complex models — effective control variates are simply unavailable (see, for instance, Duffie and Glynn (1995)). As noted in Glasserman (2013b),“some exotic payoffs are so irregular that finding an effective control variate becomes a significant challenge in itself.”
Our work seeks to bridge these two approaches, offering a hybrid framework that leverages the best of both worlds. We propose a novel control variate (CV)-based methodology, termed Prediction Enhanced Monte Carlo (PEMC), which extends the classical MC framework by incorporating modern ML techniques while preserving the unbiasedness and rigorous error quantification of MC. In the context of option pricing, unlike traditional ML applications in pricing derivatives, where models are trained to directly predict option prices, we train a certain conditional expectation as CVs, where we break away from two major characteristics of traditional CVs:
  • Known mean requirement: Unlike traditional CVs that demand the auxiliary function have a known mean (a challenging condition for complex or irregular payoffs), our approach does not rely on this assumption. Instead, PEMC utilizes ML-optimized predictors that are simple to estimate and still provide effective variance reduction.
  • Per-run variance reduction: While traditional CV techniques are designed to reduce variance on a per-sample basis—pairing each sample with a corresponding control variate—PEMC takes a broader approach. It focuses on achieving variance reduction across the entire simulation scheme, rather than targeting individual sample variance reduction.
Through a careful combination of classical MC principles and ML innovation, our approach eliminates bias from the black-box nature of ML models while also enjoying the unbiasedness and variance reduction provided by the traditional CV methods. By combining these elements, PEMC provides a principled and scalable framework that addresses the shortcomings of both classical CVs and ML-based methods, achieving reliable and interpretable results even in high-dimensional or irregular settings.
1.1 Our Contribution: A Modernized View of Control Variates
We introduce Prediction-Enhanced Monte Carlo (PEMC), a novel framework for evaluating Monte Carlo, which can be viewed as a modernized view of control variates: rather than searching for a single closed-form control variate with a known mean, we train a flexible predictor that acts as a control variate over the entire scheme. Crucially, the construction maintains unbiasedness and allows for valid statistical inference—just as in MC—while leveraging the representational power of ML models to approximate conditional expectations that can be evaluated efficiently. The key innovation of PEMC lies in relaxing the classical requirement that control variates must have a known mean. This is achieved through a two-stage estimation process: the first stage uses a small number of full simulations to estimate the difference between the target and the predictor, while the second stage computes the predictor average using a large number of computationally inexpensive simulations.
This design makes the full simulation cost modest, despite harnessing a large number of cheap samples to drive down variance of the predictor. Consequently, we achieve a form of variance reduction not just on a per-path basis (as the traditional CV), but scheme-wide: the entire simulation procedure, not a given sample design, is optimized to reduce the effective scheme variance vs scheme cost tradeoff, not the per sample variance vs per sample cost tradeoff. This differs from the defining characteristic of traditional CV methods where one control variate is applied per run. Our PEMC approach modernizes this idea, offering a “scheme-wide” variance reduction that can adapt to complex models and broad ranges of parameters.
Compared to purely machine learning-based methods, PEMC transforms the role of ML in Monte Carlo tasks. Rather than relying on ML as a direct surrogate predictor for the quantity of interest—which often introduces biased approximations with unquantifiable errors—PEMC uses ML to construct a robust, unbiased estimator augmented by modernized control variates. This approach enables fast predictions while maintaining Monte Carlo as the de facto standard of evaluation.
1.2 Literature Review
The literature on enhancing MC with ML approximators is extensive, but existing ML-related methods for control variates and variance reduction often impose stringent conditions that limit their practicality. Techniques such as the reproducing Stein kernel approach (Oates et al. (2017), Li and Zhang (2023)), regularized least squares for CV construction (Portier and Segers (2018), South et al. (2022), Leluc et al. (2021)), and adaptive CV schemes (Henderson and Glynn (2002), Henderson and Simon (2004), Kim and Henderson (2007)), as well as $\mathscr{L}^{2}$ function approximation frameworks (Maire (2003)), can in theory deliver powerful improvements—including even “supercanonical” convergence rates. But these methods usually require strict assumptions to ensure that the control variate’s mean is known and that it can be easily integrated into the simulation. In many practical settings, these assumptions fail, thus limiting the application scope of these methods. Even though some approaches can theoretically achieve supercanonical convergence rates, their applicability is limited by these constraints.
@@ 1.2 Literature Review
In contrast, the Prediction-Enhanced Monte Carlo (PEMC) framework overcomes these constraints by reimagining the control variate paradigm to function effectively in scenarios where auxiliary means are unknown. Unlike the aforementioned methods, PEMC aligns more closely with the recently proposed Prediction-Powered Inference (PPI) approach (Angelopoulos et al. (2023)), particularly in its hybrid methodology that integrates valid, unbiased and classical statistical principles with modern machine learning (ML) techniques. This aligns with recent trends in the evaluation of generative models Zrnic and Candès (2024), such as large language models (LLMs), where fast and interpretable metrics are increasingly critical. In this context, PPI-type of approaches highlight the value of combining human annotations and automatic predictions to yield unbiased estimates of model performance, even in low-label settings. This innovation is particularly valuable in high-dimensional or irregular settings, such as LLM evaluation, where annotated data may be sparse, but unannotated or simulation-based samples are plentiful Boyeau et al. (2024), Eyre and Madras (2024).
@@
Traditional variance reduction and numerical approximation techniques have a long and established history in computational finance and stochastic modeling. One example conceptually related to PEMC is the Multilevel Monte Carlo (MLMC) method, as introduced and popularized by Giles (2015). MLMC exploits a hierarchy of discretizations (e.g., finer and coarser time steps in an SDE simulation) to reduce variance at a fraction of the cost of a naive Monte Carlo approach. Its basic premise involves coupling simulations at different resolution levels so that differences between levels have lower variance, ultimately leading to significant computational savings. Despite its solid theoretical foundations, MLMC can be challenging to implement in practice: constructing suitable couplings between the fine and coarse approximations, ensuring stability, and tuning parameters to achieve the promised complexity gains are all nontrivial tasks. However, adopting a coupling perspective provides valuable insights into how PEMC manages the relationship between the target and the predictor.
The advent of machine learning (ML) has spurred a wave of research aiming to accelerate and improve financial computation, including option pricing. Techniques such as neural network surrogates can approximate pricing functions or Greeks, drastically cutting down on runtime (Hutchinson et al. (1994), Buehler et al. (2019)). However, a major drawback is the lack of rigorous error controls and biases introduced during training.
Our paper is also related to the recent stream of literature in causal inference, where methods aim to mitigate challenges suc as selection bias and confounding. Doubly robust estimators serve as a key example (Bang and Robins (2005)). These estimators maintain consistency even if only one of two specified models—either the outcome model or the propensity score model—is correctly specified. Their inherent idea of robustness through additional predictor resonates with the philosophy behind control variates, as well as PEMC, where pairing a complex estimator with a suitably chosen secondary construct can yield improved stability and reduced variance.
@@
The rest of the paper is organized as follows. In Section 2, we provide a comprehensive background on PEMC. In Section 3, we provide a detailed and comprehensive analysis of the PEMC estimator, presenting both theoretical results and practical illustrations. In Section 4, we apply PEMC it to real-world exotic option pricing problems of variance swaps pricing under stochastic (local) volatility models and the pricing of swaptions under the Heath-Jarrow-Morton (HJM) framework. We conclude the paper in Section 5 with discussion on broader implications and potential directions for future research.
2 Prediction-Enhanced Monte Carlo: Framework
In this section, we describe how to apply the “Prediction-Enhanced Monte Carlo” Framework for a general simulation evaluation. While this framework has broad applicability, for clarity and concreteness, we will present it in the context of options pricing. This choice of notation and presentation aligns with our subsequent examples and reflects a significant area of application. However, it’s important to note the principles and methodology we discuss here can be readily adapted to a wide range of problems beyond financial derivatives.
To effectively illustrate our proposed framework, facilitate a comprehensive comparison with traditional methods, and enhance the reader’s understanding, we employ Asian option pricing as a running example throughout in this section. Asian options are path-dependent derivatives whose payoffs depend on the average price of the underlying asset over a specified period. As mentioned in Boyle et al. (1997) and Dingec et al. (2015), closed-form solutions for the pricing of arithmetic average Asian options are generally not available, but arithmetic average Asian options serve as an exemplary case where control variate methods, based on closed-form pricing of the geometric average Asian option under well-known models, can drastically reduce variance and enhance the efficiency of Monte Carlo simulations. However, as we will demonstrate, our new framework achieves the same level of variance reduction without requiring explicit knowledge of the construction of a control variate or its closed-form mean. This significantly expands the applicability and effectiveness of control variate techniques in arbitrarily complex models.
2.1 Problem Setup

### v2

@@
Abstract.
For many complex simulation tasks spanning areas such as healthcare, engineering, and finance, Monte Carlo (MC) methods are invaluable due to their unbiased estimates and precise error quantification. Nevertheless, Monte Carlo simulations often become computationally prohibitive, especially for nested, multi-level, or path-dependent evaluations lacking effective variance reduction techniques. While machine learning (ML) surrogates appear as natural alternatives, naïve replacements typically introduce unquantifiable biases. We address this challenge by introducing Prediction-Enhanced Monte Carlo (PEMC), a framework that leverages modern ML models as learned predictors, using cheap and parallelizable simulation as features, to output unbiased evaluation with reduced variance and runtime. PEMC can also be viewed as a “modernized" view of control variates, where we consider the overall computation-cost-aware variance reduction instead of per-replication reduction, while bypassing the closed-form mean function requirement and maintaining the advantageous unbiasedness and uncertainty quantifiability of Monte Carlo.
We illustrate PEMC’s broad efficacy and versatility through three examples: first, equity derivatives such as variance swaps under stochastic local volatility models; second, interest rate derivatives such as swaption pricing under the Heath-Jarrow-Morton (HJM) interest-rate model.Finally, we showcase PEMC in a socially significant context—ambulance dispatch and hospital load balancing—where accurate mortality rate estimates are key for ethically sensitive decision-making.
Across these diverse scenarios, PEMC consistently reduces variance while preserving unbiasedness, highlighting its potential as a powerful enhancement to standard Monte Carlo baselines.
Equal Contribution. 1Morgan Stanley. 2Department of IEOR, Columbia University. The Work was done during the internship at Morgan Stanley in the Machine Learning Research group.
     Corresponding author: fengpei.li@morganstanley.com
@@ Abstract.
Li et al.
Key words: Monte Carlo methods, Exotic Options Pricing, Control Variates, Variance Swaps, Swaptions.
1. Introduction
Monte Carlo (MC) is a fundamental tool for evaluating complex stochastic models where closed-form solutions are unavailable. Its main strengths—unbiasedness and rigorous uncertainty quantification—make it essential in high-stakes, high-complexity applications like healthcare resource allocation and financial risk management, where fairness or regulatory compliance are crucial. For example, transparent and bias-aware methods in healthcare are vital to ensure equitable outcomes and prevent disparities in care access [110, 76, 41]. Similarly, Monte Carlo estimation is widely used in financial risk management to implement hedging strategies, rebalance portfolios, and establish reliable error bounds, as even small biases in financial simulations can accumulate over time and lead to significant errors [30].
Despite these advantages, classical MC faces significant challenges in nested, multilevel, or path-dependent simulations that are difficult to parallelize. Complex stochastic dynamics can make each sample path extremely costly to generate—sometimes taking hours for one sample (e.g., credit valuation adjustment [62])—and effective variance reduction methods, such as control variates, are often unavailable [9]. Consequently, MC’s inherent $\mathcal{O}(1/\sqrt{n})$ convergence rate often prohibits real-time decision-making with adequate precision. This is also problematic in exotic option pricing, where sample paths are generated sequentially under complex dynamics, making parallelization unavailable [60]. The issue is further exacerbated when many evaluation tasks need to be repeatedly conducted in a timely manner. In application domains such as option pricing in computational finance, such a computational barrier has been pointed out prominently, e.g., [73, Chapter 21] or [30, Chapter 4].
@@ 1. Introduction
Conversely, machine learning (ML) based predictive modeling has emerged as a fast alternative for approximating MC evaluations, including pricing financial derivatives [74, 52, 15]. Neural networks and other ML techniques can learn complex mappings and, once trained, can generate predictions at negligible marginal cost - what practitioners often call “fast evaluation" [72]. However, the black-box nature and the lack of inherent statistical control lead to unquantifiable biases, rendering them unsuitable for tasks requiring rigorous error quantification, e.g., high-stakes applications in risk management [74, 32]. Thus, MC and ML represent opposite ends of the spectrum: MC is reliable and uncertainty quantifiable but intrinsically slow, while ML is computationally fast but falls short in statistical guarantees.
Given this trade-off between the speed of ML and the reliability of MC, a natural question arises: can we design a method that combines both advantages while avoiding their respective pitfalls? In this paper, we answer affirmatively by introducing the Prediction-Enhanced Monte Carlo (PEMC) framework. PEMC leverages predictive ML modeling to replace a large portion of costly simulations with cheap/parallelizable samples from ML models, reducing variance on a scheme-wide level. Moreover, despite relying on black-box predictive models, PEMC maintains unbiasedness, balancing computational savings with rigorous statistical guarantees.
1.1. Contribution
The PEMC innovates in two aspects. First, PEMC integrates ML predictive models into any MC baseline while preserving unbiasedness and error quantification of MC. Second, we show how the predictive modeling of ML combined with cheap, parallelizable samples will result in a scheme-wide reduction of variance and costly samples, enhancing real-time efficiency for complex MC simulations. PEMC method can also be viewed as a modernized variant of the control variate (CV) technique, a classical approach that reduces Monte Carlo variance via auxiliary outputs whose known mean offsets random fluctuations [60, 9]. Unfortunately, the classical CV requirement of a closed-form mean of the a correlated auxiliary output easily becomes prohibitive in many modern applications or complicated stochastic systems. In such cases, there simply is no readily available analytic expression for the auxiliary variable’s expectation, effectively barring the use of a CV.
The PEMC framework bridges the above gap and substantially enlarge the scope of CV to many ML candidates. The key idea is the following: In the traditional CV framework, the known mean of CV is used to achieve a per-replication variance reduction while maintaining unbiasedness, i.e., the variance is reduced even if we run just one simulation replication. This is a powerful ability, but as mentioned above it is achieved at the expense of severely narrowing the applicability of CV by requiring closed-form mean function. However, if the CV itself is computationally cheap in the sense of being estimatable with fast simulation, then we can afford to run plenty of additional MC separately from the original MC to estimate its mean. This would preserve unbiasedness and overall estimation variance relative to the total computation cost, we could achieve a scheme-wide efficiency improvement. That is, in the PEMC framework, we consider scheme-wide instead of per-replication variance reduction, through which we relax the crucial requirement of known CV mean, and this in turn allows us to substantially expand the CV methodology to cover modern ML models. The alignment of our framework with the use of ML predictions prompts us to regard PEMC as a modernized view of CV.
We demonstrate PEMC’s efficacy across multiple MC applications. First, we consider two exotic financial derivative-pricing scenarios: (i) variance swap pricing under a stochastic local volatility model, a fundamental instrument for volatility trading due to its payout on realized variance; and (ii) swaption pricing within the Heath-Jarrow-Morton (HJM) interest rate framework, a sophisticated yet widely used derivative by financial institutions for hedging and speculation. Both instruments are highly liquid, exchange-traded products but pose significant computational challenges due to the infeasibility of traditional variance reduction techniques, as they lack analytic or easily computable expected values. Additionally, we illustrate PEMC’s broad applicability through an ambulance dispatch and hospital load-balancing simulation, where unbiased mortality rate estimates critically inform ethical and efficient resource allocation. These examples collectively highlight PEMC’s capability to leverage machine learning predictions to effectively manage computational complexity while maintaining unbiased estimation and reliable error quantification.
1.2. Literature Review
The literature on enhancing Monte Carlo (MC) simulations with machine learning (ML) approximators is extensive, yet many existing ML-based methods for control variates (CV) and variance reduction impose restrictive conditions that limit their practical applicability. Approaches such as reproducing Stein kernels [100, 88], Neural CV derived from Stein’s identity [129, 82], regularized least squares for CV construction [106, 121, 85], adaptive CV schemes [66, 67, 79], and $\mathscr{L}^{2}$ function approximation frameworks [93] have demonstrated theoretical promise with potentially rapid convergence rates. Recent theoretical work by [18] provides important insights into when regression-adjusted control variates can achieve optimal performance, showing that their effectiveness depends critically on function smoothness. However, these methods typically require knowledge of the control variate mean or impose specific structural assumptions that can be difficult to check. As a result, while these existing methods offer strong theoretical foundations, their applicability in complex financial modeling scenarios can be sometimes limited. Multilevel Monte Carlo (MLMC) [58], another prominent variance reduction approach, has shown effectiveness through coupling simulations at different discretization levels but is often challenging to implement, requiring careful design for stable coupling [107, 19, 112, 57, 115]. Our Prediction-Enhanced Monte Carlo (PEMC) framework addresses these limitations comprehensively by eliminating the necessity of known CV means, enabling the direct use of modern ML predictors as control variates, while rigorously preserving unbiasedness.
In a different line of literature, the work of Emsermann and Simon [49] on quasi control variates and the work of [102] on control variates using Estimated Means (CVEMs) share notable similarities with our approach, as it relies on an estimated mean (rather than a priori known, and thus not a traditional control variate) of a quantity that is expected to exhibit a strong correlation with the random variable of interest. However, our methodology distinguishes itself in several critical ways. First, our primary motivation centers on integrating modern ML capabilities to train a suitable control variate, rather than relying on a fixed quantity observed to be correlated with the target (which may not always be obvious or even available, especially for our options pricing examples). In other words, we can create an effective control variate from an easily estimable quantity, even if its naive form does not exhibit strong correlation with the target. This approach circumvents the challenges of identifying a closed-form mean and strictly relying on high correlation-based pairing. Second, instead of adopting an adaptive algorithm as they do, we focus on a pre-training approach, demonstrating that our performance advantages would not emerge under their adaptive scheme. Third, our analysis is not anchored exclusively in asymptotic arguments. We provide a formulation that can be interpreted in both finite-sample and asymptotic regimes, offering more direct and flexible theoretical guarantees.
@@ 1.2. Literature Review
In contrast, the Prediction-Enhanced Monte Carlo (PEMC) framework overcomes these constraints by reimagining the control variate paradigm to function effectively even when auxiliary means are unknown. Unlike previously discussed methods, PEMC aligns closely with the recently proposed Prediction-Powered Inference (PPI) approach [6], particularly in its hybrid methodology that integrates classical, unbiased statistical principles with modern machine learning (ML) techniques. This hybrid strategy is becoming increasingly important in evaluating generative models such as large language models (LLMs), where rapid, interpretable, and unbiased metrics are essential [135, 22, 51]. PPI-type approaches combine sparse human annotations with abundant ML-generated labels, facilitating accurate model performance evaluation even in low-label scenarios.
Additionally, these methodologies intersect with a rich literature in causal inference, specifically doubly robust estimators [10, 55, 113] and orthogonal statistical learning [36, 54]. Doubly robust methods construct additional robustness layers by pairing sophisticated predictive models with auxiliary estimators, ensuring that minor errors in one model component do not significantly bias the overall estimation. This philosophy closely mirrors the logic of control variates, and by extension, the core principles of our PEMC framework.
@@
Traditional variance reduction and numerical approximation techniques also have a long and established history in computational finance and stochastic modeling. One example conceptually related to PEMC is the Multilevel Monte Carlo (MLMC) method, as introduced and popularized by [58]. MLMC exploits a hierarchy of discretizations (e.g., finer and coarser time steps in an SDE simulation) to reduce variance at a fraction of the cost of a naive Monte Carlo approach. Its basic premise involves coupling simulations at different resolution levels so that differences between levels have lower variance, ultimately leading to significant computational savings. Despite its solid theoretical foundations, MLMC can be challenging to implement in practice: constructing suitable couplings between the fine and coarse approximations, ensuring stability, and tuning parameters to achieve the promised complexity gains are all nontrivial tasks. Nevertheless, the class of MLMC methods and their adoption of a coupling perspective provide extremely valuable insights into how PEMC manages the relationship between the target and the predictor [107, 19, 112, 57, 115].
In computational finance, the emergence of machine learning (ML) has also sparked significant research efforts aimed at enhancing the speed and accuracy of financial computations, including option pricing. By using techniques such as neural network surrogates, these methods can approximate pricing functions or the Greeks at a fraction of the computational cost of traditional numerical schemes [74, 32, 119, 114, 84]. Early work by hutchinson1994nonparametric [74], for instance, used neural networks to relax stringent modeling assumptions, while more recent approaches [32, 119, 114, 84] demonstrate that deep architectures can effectively handle high-dimensional and path-dependent pricing problems. However, these purely data-driven approximations typically lack the rigorous error controls afforded by more classical methods. As a result, biases introduced during training and potential overfitting can lead to systematic mispricings, particularly under market regimes not well-represented in the training data. Such issues underscore the need for frameworks that combine the computational advantages of ML with robust theoretical guarantees to ensure reliable and unbiased pricing outcomes. On the other hand, within option pricing, even though both MC and PDE-based methods are widely used, it is known that PDE methods generally cannot be used to price options (i.e., payoff depends on the price path). While special cases allow for PDE formulations through state-space augmentation or boundary condition adjustments, these remain as exceptions rather than the rule [116].
@@
The rest of the paper is organized as follows. In Section 3, we provide a comprehensive background on PEMC. In Section 4, we provide a detailed and comprehensive analysis of the PEMC estimator, presenting both theoretical results and practical illustrations. In Section 5, we apply PEMC it to real-world exotic option pricing problems of variance swaps pricing under stochastic (local) volatility models and the pricing of swaptions under the HJM framework. We conclude the paper in Section 6 with discussion on broader implications and potential directions for future research.
2. Prediction-Enhanced Monte Carlo
We provide an overview and motivational explanation of our PEMC framework. Suppose we want to produce an unbiased evaluation of the quantity
$\mu(\boldsymbol{\theta})=\mathbb{E}_{\boldsymbol{\theta}}[f_{\boldsymbol{\theta}}(\boldsymbol{Y})]$ (1)
for arbitrary, potentially high-dimensional parameter $\boldsymbol{\theta}\in\boldsymbol{\Theta}\subseteq\mathbb{R}^{d_{\theta}}$, where $\Theta$ is the parameter space. Here, $\boldsymbol{Y}\in\mathbb{R}^{d_{y}}$ represents a high-dimensional random vector governed by $\mathbb{E}_{\boldsymbol{\theta}}$ and $f_{\boldsymbol{\theta}}$ is the evaluation function $\mathbb{R}^{d_{y}}\to\mathbb{R}$. In general, both $f_{\boldsymbol{\theta}}$ and $\mathbb{E}_{\boldsymbol{\theta}}$ can depend on $\boldsymbol{\theta}$. The goal is to produce an unbiased estimate of $\mu(\boldsymbol{\theta})$ for any $\boldsymbol{\theta}\in\boldsymbol{\Theta}$, while the primary bottleneck stems from costly and time-consuming simulation of $\boldsymbol{Y}$.
When $\boldsymbol{Y}$ can be simulated and $f$ evaluated, then standard MC amounts to generating many $f(\boldsymbol{Y}_{i})$, $i=1,\ldots,n$ and outputting their average. While this approach is demanding when $f$ and $\boldsymbol{Y}$ are expensive, it is unbiased and its statistical error can be quantified straightforwardly via standard confidence intervals.
To reduce computation, one can approximate $\mu(\boldsymbol{\theta})$ with an ML regression model, say $\hat{\mu}(\boldsymbol{\theta})$, which can be evaluated more cheaply than simulating many $f_{\boldsymbol{\theta}}(\boldsymbol{Y}_{i})$’s. This natural approach, however, would face challenges at multiple levels when reliability and error quantification is important: From a theoretical standpoint, ML generalization bounds typically assess training errors only within the model class, i.e., $\hat{\mu}(\boldsymbol{\theta})-\mu^{*}(\boldsymbol{\theta})$ where $\mu^{*}(\cdot)$ is the best in-class model that can be different from $\mu(\cdot)$. Moreover, even assuming $\mu^{*}(\cdot)$ is a “rich" class, these bounds still only assess errors in an aggregate, worst-case fashion, i.e., high-probability bounds treating $\boldsymbol{\theta}$ as a random covariate. On the practical front, training-testing split or cross-validation succumbs to similar issues, in that it cannot measure bias against $\mu(\cdot)$, nor at the level of each $\boldsymbol{\theta}$. That is, in terms of reliability and uncertainty handling, MC appears superior by providing accurate, straightforward and instance-specific error assessment on each estimate of $\mu(\boldsymbol{\theta})$. None of these advantages are available for ML despite its computational edge.
An idea to leverage the respective advantages of MC and ML is to use control variate (CV), a classical variance reduction method. CV aims to reduce the error of each MC run, consequently the required replication size, by utilizing information from “auxiliary” simulation outputs. Specifically, in generating $f(\boldsymbol{Y}_{i})$, suppose we can also generate auxiliary output $\boldsymbol{X}_{i}$ and evaluate $g(\boldsymbol{X}_{i})$ for some function $g$. Then we can use $f(\boldsymbol{Y}_{i})-g(\boldsymbol{X}_{i})+\mathbb{E}[g(\boldsymbol{X})]$ as an unbiased replicate for $\mathbb{E}[f(\boldsymbol{Y})]$. In particular, if $f(\boldsymbol{Y})$ and $g(\boldsymbol{X})$ are highly correlated, this replicate will have a smaller variance than $f(\boldsymbol{Y}_{i})$.
Naturally, one can think of using ML to train a good predictor $g(\boldsymbol{X})$ for $f(\boldsymbol{Y})$, which would be highly correlated with $f(\boldsymbol{Y})$ and serve as an effective CV. However, taking aside the question of how to construct $\boldsymbol{X}$ and $g$, the CV method requires knowledge of $\mathbb{E}[g(\boldsymbol{X})]$, the closed-from mean on the auxiliary output. This knowledge is the key convertible into variance reduction. Yet, it is seldom available and has apparently severely limited the historical applicability of CV [61] and arguably even the more recent ML-inspired approaches [129, 121, 88].
We introduce Prediction-Enhanced Monte Carlo (PEMC), a framework that fulfills these objectives. PEMC leverages the classical control variate idea but generalizes its variance-reduction notion to a scheme-wide, modernized view, allowing for effective integration of MC with modern ML. Concretely, PEMC aims to estimate (1) by
$PEMC(\boldsymbol{\theta}):=\frac{1}{n}\sum_{i=1}^{n}\big{(}f_{\boldsymbol{\theta}}(\boldsymbol{Y}_{i})-g(\boldsymbol{\theta},\boldsymbol{X}_{i})\big{)}+\frac{1}{N}\sum_{j=1}^{N}g(\boldsymbol{\theta},\boldsymbol{\tilde{X}}_{j}).$ (2)
Here, $\boldsymbol{X}$ is an auxiliary variable, coupled with $\boldsymbol{Y}$, that serves as features for predicting $f(\boldsymbol{Y})$ and is also chosen to have a marginal distribution that is both efficiently simulatable and highly parallelizable, allowing large number $N$ independent samples of $\tilde{\boldsymbol{X}}$ with the same marginal as $\boldsymbol{X}$ to be generated at low cost. The function $g$ is a prediction model trained via squared-error loss using MC sample under $(\boldsymbol{X},\boldsymbol{Y})\sim\mathbb{P}_{\boldsymbol{\theta}}$ for various $\boldsymbol{\theta}\sim\boldsymbol{\Theta}$, which aims to approximate the function $g\bigl{(}\boldsymbol{\theta},\boldsymbol{X}\bigr{)}\approx\mathbb{E}_{\boldsymbol{\theta}}\bigl{[}f_{\boldsymbol{\theta}}\bigl{(}\boldsymbol{Y}\bigr{)}\big{|}\boldsymbol{\theta},\boldsymbol{X}\bigr{]}$. As we will see, the selection of an appropriate $\boldsymbol{X}$ is a pivotal component of PEMC, analogous to feature selection in common ML tasks. The key lies in selecting $\boldsymbol{X}$ to be both a feature for predicting $f(\boldsymbol{Y})$ and an efficiently simulatable, parallelizable random variable. Finding a good $\boldsymbol{X}$ often requires domain knowledge. In particular, given the whole path $\boldsymbol{Y}$, one can choose $\boldsymbol{X}:=\phi(\boldsymbol{Y})$ to be some low-dimensional transformation $\phi$ of $\boldsymbol{Y}$ (e.g., ingredients for constructing $\boldsymbol{Y}$). For instance, for SDEs, one can define $\boldsymbol{X}$ as the sum of driving Brownian increments that generate $\boldsymbol{Y}$. This choice could capture a substantial portion of the SDE’s variance, making $\boldsymbol{X}$ a good feature for predicting $f(\boldsymbol{Y})$, while its Gaussianity allows it to be generated cheaply and in parallel. We explore this in detail in subsequent sections.
Many implementation details remain unspecified in the basic description above. For instance, how exactly to train g, how to select the sample sizes $n$ and $N$, and how to adapt PEMC to a conditional setting where $\theta\mapsto\boldsymbol{X}(\theta)$, $\boldsymbol{Y}(\theta)$, and $\mathbb{E}_{\theta}$ all depend on some parameter $\theta$. The next sections will address these questions in depth, providing both theoretical justification and practical guidelines for deploying PEMC in diverse scenarios.
3. Prediction-Enhanced Monte Carlo: Framework via Example
In this section, we employ Asian option pricing as a running example on how to apply PEMC. Asian options are path-dependent derivatives whose payoffs depend on the average price of the underlying asset over a specified period. As mentioned in [23] and [44], closed-form solutions for the pricing of arithmetic average Asian options are generally not available.
3.1. Problem Setup